OpenStudy (anonymous):

Can anyone help me with this? Find the derivative: f(x)=ln(sin(e^x))+(1/2(tan))[(ln(x+1))]

OpenStudy (anonymous):

It's going to be a pain in the retrice but sure. Break that pellet into two derivatives. Let's do this pellet. 1) ln(sin(e^x)) derived is [ 1 / ( sin(e^x)) ] * [ sin(e^x) ] ' = [ 1 / ( sin(e^x)) ] * [ cos(e^x) *e^x] = ctg (e^x) * e^x 2) (1/2(tan))[(ln(x+1)) derived is ... wait, did you mean 1/2 * tan ( ln(x+1) ) by any chance ?

OpenStudy (anonymous):

1/2tan[ln(x+1)]

OpenStudy (amorfide):

\[f(x)=\ln(\sin(e^{x})) \] let u=sin(e^x) to get f(x)=lnu \[f'(x)= \frac{ f'(u) }{ f(u) }\] \[f'(x)=\frac{ e^{x}\cos(e^{x}) }{ \sin(e^{x}) }\] \[\frac{ cosx }{ sinx }=cotx\] \[f'(x)=e^{x}\cot(e^{x})\] now another function \[g(x)=\frac{ 1 }{ 2 }\tan(\ln(x+1))\] \[g'(x)=\frac{ \sec^{2}\ln(x+1) }{ 2(x+1)}\] add them together to get the differential \[f'(x)=e^{x}\cot(e^{x}) + \frac{ \sec^{x}\ln(x+1) }{ 2(x+1) }\]

OpenStudy (amorfide):

to differentiate the tan function I let ln(x+1)=u so I have \[g(x)=\frac{ 1 }{ 2 }\tan(u)\] \[g'(x)=\frac{ 1 }{ 2 }\sec^{2}u \] then I multiply by the differential of u

OpenStudy (amorfide):

oops that should be u in the brackets g(u) g'(u)

OpenStudy (amorfide):

hope that helps

OpenStudy (anonymous):

wait so what's the answer

OpenStudy (anonymous):

i think i got something different

OpenStudy (amorfide):

hope that helps

OpenStudy (amorfide):

I gave you the answer

OpenStudy (amorfide):

http://prntscr.com/6hb0bp

OpenStudy (anonymous):

oops didn't see that my chat is lagging for some reason

OpenStudy (amorfide):

need me to explain anything?

OpenStudy (anonymous):

Yes I am having trouble with the trapezoid and simpsons rule i was wondering if you could help me on that

OpenStudy (amorfide):

I meant on this question lol post your next question in a new question and I will try to help

OpenStudy (anonymous):

Ok I will post a new question.